OpenStudy (anonymous):

Prove by induction:

OpenStudy (anonymous):

\[\sum_{i=1}^{n} i = n(n+1)/2\]

OpenStudy (anonymous):

i have up to the inductive hypothesis is \[\sum_{i=1}^{k} i = k(k+1)/2 = (k^2+k)/2\]

OpenStudy (anonymous):

What I have so far for the next part is: \[\sum_{i=1}^{k+1} i = (k+1)(k+2)/2 = \frac{ k^2+2 }{ 2 }+\frac{ 2k+1 }{ 2 }\]

OpenStudy (anonymous):

I'm just not entirely sure where to go from here

ganeshie8 (ganeshie8):

you're mostly done

ganeshie8 (ganeshie8):

except for a typo

ganeshie8 (ganeshie8):

\[\large \sum_{i=1}^{k+1} i = (k+1)(k+2)/2 = \frac{ k^2+k}{ 2 }+\frac{ 2k+2 }{ 2 } \]

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

\[\large \begin{align}\\ \sum_{i=1}^{k+1} i = (k+1)(k+2)/2 &= \frac{ k^2+k}{ 2 }+\frac{ 2k+2 }{ 2 } \\~\\&= \dfrac{k(k+1)}{2} + (k+1)\end{align}\] QED

OpenStudy (anonymous):

Okay. I looked at multiple lines when I was typing my equations. However, I did mess up the 2 for whatever reason. That helps a lot. Thank you again!

ganeshie8 (ganeshie8):

You could also try this directly from induction hypothesis : \[\large \sum_{i=1}^{k} i + (k+1) = \dfrac{k(k+1)}{2} + (k+1) = \dfrac{(k+1) (k+2)}{2} \] i think thats what water is typing right now :)

OpenStudy (anonymous):

He he he.. I was typing my next question.. :P

ganeshie8 (ganeshie8):

lol i have no clue wish i could attempt these .-.

OpenStudy (anonymous):

Sometimes, I need some space for LATEX works in the question.. :P

OpenStudy (anonymous):

So to end these kinds of problems, is it mostly a matter of getting the induction hypothesis plus something else that proves the question? I'm still kindof new at induction and trying to figure it out.

ganeshie8 (ganeshie8):

induction is a beautiful way of proving a statement is valid for ALL integers, its not for ending prolems :P

ganeshie8 (ganeshie8):

you know how/why induction proof works right ? have u seen a proof proving induction method ?

OpenStudy (anonymous):

In Induction, there are just three steps you have to follow to prove anything you are asked to.. Checking the validity at n = 1. Secondly, assuming that for n = k, it is true. Thirdly, for n = K+1, prove that it is true..

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

I have seen them, I'm still getting used to not actually solving the problems, but proving that they can be solved. I can usually get everything up to the last ending statement as you can tell. It's just a matter of not really knowing what the right ending point is

ganeshie8 (ganeshie8):

Exactly ! its always a trial and error business i like starting with "n=k" and transitioning to "n=k+1"

ganeshie8 (ganeshie8):

the other way is starting with "n=k+1" statement and plugging in "n=k" stuff which is okay but my personal preference is the former method

OpenStudy (anonymous):

for example, one of my other questions proves that n^3 +2n is divisible by 3. I have the inductive hypothesis of k^3+2k and the k+1 of (k^3+2k)+3k^2+3k+3, which is divisible by 3, but I'm not really sure what I should do with the inductive hypothesis

ganeshie8 (ganeshie8):

to be honest i also don't know need to try different things before arriving at something useful

ganeshie8 (ganeshie8):

since k^3+2k is divisible by 3, can we write \[\large k^3 + 2k = 3t\] for some \(t\) ?

ganeshie8 (ganeshie8):

k+1 step : ( `k^3+2k`)+3k^2+3k+3 here, you can replace `k^3+2k` by \(3t\) precicely because of the induction hypothesis, right ?

ganeshie8 (ganeshie8):

you cannot conclude "k+1" expression is divisible by 3 without using the induction hypothesis

OpenStudy (anonymous):

That's what it looks like I may have to do. I actually found this exact question online now. That's what it is saying to do. and that does make it work out very nicely

OpenStudy (anonymous):

Thank you again! I'm going to bed now. I have to get up soon :(

ganeshie8 (ganeshie8):

have good sleep :) go thru this proof of induction method when free http://prntscr.com/4juawh